OpenStudy (anonymous):

The area A of a circle is a function with respect to its radius r: A = pi r^2. (dA)/(dr) =

OpenStudy (nenadmatematika):

dA=2pi*r

OpenStudy (anonymous):

that's the equation so I just need to plug in the numbers, right?

OpenStudy (nenadmatematika):

you don't have to plug in anything if this is the complete text

OpenStudy (anonymous):

what is the star between pi and r represent?

OpenStudy (turingtest):

times

OpenStudy (anonymous):

Remember that pi is just a number and the radius, r, is your independent variable. So you take the derivative using the power rule. So, A'= 2*(pi)r^(2-1), or just 2(pi)r.

OpenStudy (turingtest):

\[A=\pi r^2\]\[\frac{dA}{dt}=\frac d{dt}(\pi r^2)=\pi\frac d{dt}r^2=\pi(2r)=2\pi r\]

OpenStudy (turingtest):

right, power rule

OpenStudy (anonymous):

that what I got ok thanks

OpenStudy (nenadmatematika):

you're welcome :D